Jamie is twice as old as her brother.Four years ago,she was 4 times as old as her brother.Find their ages now.

Answer by rmnavalta(7)

Let x = age of Jamie's brother
   2x = Jamie's age               ***Jamie is twice as old as her brother

Let's use a table.
  
             |   Age NOW  |  Age 4 YEARS AGO |
Jamie        |      2x    |      2x-4        |
Her Brother  |      x     |      x-4         |

Four years ago, she (Jamie) was(is equal to) 4 times as old as her brother.
                    2x-4          =             4 times (x-4)

Set-up the equation.   2x-4 = 4(x-4)
               
Solve.         2x-4 = 4x-16       -->Simplify 4(x-4).
           2x-4 -4x = 4x-16 -4x   -->Subtract 4x on both sides.
           -2x-4 +4 = -16 +4      -->Add 4 on both sides.
         (-2x)/(-2) = (-12)/(-2)  -->Divide both sides by -2.
                  
                  x = 6 years (age of Jamie's brother
  
Jamie's age = 2x = 2(6) = 12 years

Answer: Jamie is 12 years old and her brother is 6 years old.